In 12 years time Ruth will be three times as old as she was 6 years ago . Find Ruth's age now
Answer by Timnewman(323)

Hi dear,
Let ruth's age be x
Six years ago,
Ruths age=x-6
In 12years time,
Ruth's age =x+12
But,
x+12=3(x-6)
x+12=3x-18
3x-x=12+18
2x=30
x=15
Ruth's age now is 15yrs
HOPE THIS HELPS?
::TIMNEWMAN::